sove pde d²z/dx²+d²z/dxdy+dz/dy-Z=e^-x
1
Expert's answer
2020-06-09T18:28:42-0400

"\frac{\partial^{2}z}{\partial^{2}x^{2}} + \frac{\partial^{2}z}{\partial x\partial y} + \frac{\partial z}{\partial y} -z = e^{-x}"


Equation can be written as:


"(D^{2} + DD' + D' - 1)z = e^{-x}"


"(D + 1)(D + D' -1)z = e^{-x}"

Comparing above equation with

"(D -m_1 D' - \alpha_1)(D -m_2 D' - \alpha_2) = F(x,y)"

so "m_1 = 0, m_2 = -1, \alpha_1 = -1, \alpha_2 = 2"


So C.F. of the equation is

"C.F. = e^{-x}\phi_1(y) + e^{x}\phi_2(y-x)"


then PI is calculated as

"P.I. = \frac{1}{(D+1)(D+D'-1)}e^{-x} = \frac{1}{(D+1)(D-1)}e^{-x}" as exponential has no y.


"P.I. = \frac{1}{(D+1)(D-1)}e^{-x}" "= e^{-x} \frac{1}{D(D-2)}1" "= -\frac{xe^{-x}}{2}"


So solution to the problem is

"z(x,y) = e^{-x}\phi_1(y) + e^{x}\phi_2(y-x) - \frac{xe^{-x}}{2}"


Solution can be verified by taking partial derivatives and put them in differential equation.